A=1/2 bh for h
solve the equation
the largest angle in a triangle is 3 more than 4 times as big as the smallest angle. The middle angle is 2 less than twice the smallest. What is the measure of the largest angle?

Answer by midwood_trail(310)

I'll do the equation.
Solve for h.
A = 1/2 (bh)
We can write 1/2 (bh) like bh/2 for easy reading. They both mean the same thing.
We now have A = bh/2
Multiplying both sides of the equation by 2 and we get:
2A = bh
To solve the equation for h, we want to get h alone.
To do so, divide both sides of the equation by b.
2A/b = h

Did you follow?